Complete this table of values for four aqueous solutions at 25 °C. OH pH Given Number Number Number Solution A: 6.9 x 10-9 Number Given Number Number Solution B:D 9.0x 10-10 M Number Number Given: Number Solution C: 3.92 Number Number Number Given Solution D: 2.00
Answer #1

A)

we have below equation to be used:

[OH-] = Kw/[H+]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[OH-] = (1.0*10^-14)/[H+]

[OH-] = (1.0*10^-14)/(6.9*10^-9)

[OH-] = 1.449*10^-6 M

we have below equation to be used:

pH = -log [H+]

= -log (6.9*10^-9)

= 8.1612

we have below equation to be used:

pOH = -log [OH-]

= -log (1.449*10^-6)

= 5.8388

Answers:

[H+] = 6.9*10^-9

[OH-] = 1.449*10^-6

pH = 8.1612

pOH = 5.8388

B)

we have below equation to be used:

[H+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H+] = (1.0*10^-14)/[OH-]

[H+] = (1.0*10^-14)/9.0E-10

[H+] = 1.111*10^-5

we have below equation to be used:

pH = -log [H+]

= -log (1.111*10^-5)

= 4.9542

we have below equation to be used:

pOH = -log [OH-]

= -log (9*10^-10)

= 9.0458

Answers:

[H+] = 1.111*10^-5

[OH-] = 9*10^-10

pH = 4.9542

pOH = 9.0458

C)

POH = 14 - pH

= 14 - 3.92

= 10.08

we have below equation to be used:

pH = -log [H+]

3.92 = -log [H+]

log [H+] = -3.92

[H+] = 10^(-3.92)

[H+] = 1.202*10^-4 M

we have below equation to be used:

pOH = -log [OH-]

10.08 = -log [OH-]

log [OH-] = -10.08

[OH-] = 10^(-10.08)

[OH-] = 8.318*10^-11 M

Answers:

[H+] = 1.202*10^-4

[OH-] = 8.318*10^-11

pH = 3.92

pOH = 10.08

D)

we have below equation to be used:

PH = 14 - pOH

= 14 - 2

= 12

we have below equation to be used:

pH = -log [H+]

12 = -log [H+]

log [H+] = -12

[H+] = 10^(-12)

[H+] = 1*10^-12 M

we have below equation to be used:

pOH = -log [OH-]

2 = -log [OH-]

log [OH-] = -2

[OH-] = 10^(-2)

[OH-] = 1*10^-2 M

Answers:

[H+] = 1.00*10^-12

[OH-] = 1.00*10^-2

pH = 12.00

pOH = 2.00

Feel free to comment below if you have any doubts or if this answer do not work